RC meeting and Run Plan, Monday March 8th
From SBS wiki
Return to the RC Meeting Calendar
Contents
RC Report
- RC: Wenliang (billlee) Li, billlee@jlab.org
Summary
- RC Daily report: https://logbooks.jlab.org/entry/4145213
CW could be sent to Hall A. A & C couldn't be brought up simultaneously (there could still be an issue). He3 target was not ready due to a position switch issue. Collected 4 hours of reference Cell data (H2) overnight. Overall, we have production beam onsight, and MCC day shifters are eager to give Hall A high-quality beam. We should be back online soon.
See a comprehensive Owl operation summary by Will Tireman: https://logbooks.jlab.org/entry/4145057
Recent Issues
- 4th pass separator connector replacement: https://logbooks.jlab.org/entry/4144059
- March 7 @ 1:00 pm, we might have lost polarization due to a strange incident https://logbooks.jlab.org/entry/4144460
- March 8 @ 2:00 am, accelerator issue: it was found that the only way Hall A could run at CW with the trip is to ramp Hall C to full current and slowly bring hall A online.
- See log entry from MCC: https://logbooks.jlab.org/entry/4145006
- March 8 @ 3:00 am, target position switch mismatch.
- See MCC log: https://logbooks.jlab.org/entry/4144919
- Bill Henry's diagnostics: https://logbooks.jlab.org/entry/4144919 (this need to be part of the future target installation plan)
- Target chamber door interlock.
- Problem solved: https://logbooks.jlab.org/entry/4145291
- Some reference cell data was collected, but some issues:
- SBS GEMs are turned off
- PS trigger settings were set incorrectly: (coin trigger and single trigger were both set to 1)
- Carbon hole study position not reproducible:
- Jiwan & Bill Henry made a new rule about during the target (fur NMR during optics checkout)
Current Status
- No beam, target chamber door interlock issue is been fixed
- Accelerator is up
Run Plan
- Beam restoration
- Harp Scene
- Spot check
- 40uA production (may be lower)
Data Collection and Polarization Status
- Currently 40%: https://logbooks.jlab.org/entry/4144898
- Reduced since the laser is turned off for the position switch adjustment
Other Topics/Status Updates
- March 7 @ 9:00 pm, MCC did a beautiful Harpscan which got us parallel, sigma x=0.43mm, sigma y = 0.30mm, beam https://logbooks.jlab.org/entry/4144675
- Bill Henry found some interesting about the BBCAL trigger rate, and it might be related to the raster change see log entry below
https://logbooks.jlab.org/entry/4143923
- Carlos also did some diagnostics with Grinch LED and dark rate, see below:
https://logbooks.jlab.org/entry/4144159